# SAM-CP: Marrying SAM with Composable Prompts for Versatile Segmentation

### v1

@@ C.2 More Segmentation Examples
Appendix D Broader Impacts and Safeguards
@@
Safeguards. We have not released any data or models with a high risk of misuse. The proposed model is trained on benchmark datasets such as COCO, ADE20K, and Cityscapes. These datasets are publicly available, widely used in the field of computer vision, and have undergone comprehensive safety risk assessments. In this paper, we clearly specify the sources of the datasets and code used and provide appropriate links in the references section. Upon completion of the review process, we will make the code and data of this work publicly available to the community. This research does not involve any human subjects experiments or studies, nor does it involve crowdsourced experiments or human subjects research. All experiments were conducted using code and GPU servers.
Figure 9: More visualization of a qualitative study on how SAM-CP work .

### v2

@@ D.3 More Segmentation Examples
Appendix E Broader Impacts and Safeguards
@@
Safeguards. We have not released any data or models with a high risk of misuse. The proposed model is trained on benchmark datasets such as COCO, ADE20K, and Cityscapes. These datasets are publicly available, widely used in the field of computer vision, and have undergone comprehensive safety risk assessments. In this paper, we clearly specify the sources of the datasets and code used and provide appropriate links in the references section. Upon completion of the review process, we will make the code and data of this work publicly available to the community. This research does not involve any human subjects experiments or studies, nor does it involve crowdsourced experiments or human subjects research. All experiments were conducted using codes and GPU servers.
Figure 12: The visualization of camouflage object segmentation on COD10K to show the out-of-domain generalization. Left is the GTs and right is the results.
Figure 13: More visualization of a qualitative study on how SAM-CP works.